OpenStudy (anonymous):

Calculate the derivative of the function. f(x) = square root of 3x + x^2

OpenStudy (agreene):

\[f(x) = \sqrt{x + x^2}\] Is this right?

OpenStudy (anonymous):

no i entered it into webassign and it says incorrect

OpenStudy (agreene):

I was asking if that was the question.

OpenStudy (anonymous):

oh no, its the square root of 3x+x^2 so \[\sqrt{3x+x^2}\]

OpenStudy (agreene):

Yeah, thats what I meant, forgot that silly 3 :P \[f'(x)=\frac{2x+3}{2\sqrt{x(x+3)}}\]

OpenStudy (anonymous):

Can you show the steps or at least how you got the answer? The square root just confuses me

OpenStudy (agreene):

Use the modified chain rule: \[\frac{d\sqrt u}{du}\frac{du}{dx}\] with: u= x^2+3x d(u^(1/2))/du = 1/(2(u^(1/2))) We then have: \[\frac{\frac{d}{dx}(x^2+3x)}{2\sqrt{x^2+3x}}\] \[=\frac{2x+3}{2\sqrt{x(x+3)}}\]

OpenStudy (anonymous):

Oh I see. Thank you!